# MASQUE: A Text-Guided Diffusion-Based Framework for Localized and Customized Adversarial Makeup

### v1

@@ 5.1 Experimental Settings
Dataset. To reflect the high-quality nature of online facial images, we conduct experiments using $300$ images from each of the CelebA-HQ [14] and VGG-Face2-HQ [3] datasets, all at a resolution of $1024\times 1024$. We randomly sample $100$ identities from per dataset, each with three images: a probe image to be protected, a reference image for guidance, and a the gallery image, which is assumed to be stored in the facial recognition system for comparison.
@@
Configuration. We compare our methods against the following baselines: TIP-IM [37], the SOTA noised-based AFR, and other recent generative-based AFR methods, including AMT-GAN [12], Clip2Protect [26], and DiffAM [30]. For Clip2Protect, we denote it as C2P for simplicity, where both the impersonating version (denoted as C2P (I)) and the dodging version (i.e., C2P (D)) are assessed. We compare the performance of MASQUE with existing AFR techniques on four public FR models: IR152 [6], IRSE50 [11], FaceNet [25], and MobileFace [2]. We also evaluate AFR methods against commercial FR APIs: Face++  [^1] https://www.faceplusplus.com/face-comparing/ and Luxand  [^2] https://luxand.cloud/face-api.
@@
Figure 3: Comparison of concentration and hue accuracy across different methods for three makeup tasks: red lipstick, blue eyebrow, and pink eyeshadow. Concentration values are normalized relative to MASQUE.
For reference-based methods, we selected images from the makeup dataset [16] used during their pre-training, that best matched the given prompt. For the blue eyebrow style—introduced to assess performance on an uncommon makeup—finding a suitable reference was challenging, underscoring the limitations of reference-based methods for rare styles. To address this, we inpainted blue eyebrows on a non-makeup image from the dataset using our framework, excluding pairwise adversarial guidance, demonstrating its capacity for targeted edits without an external reference.
Evaluation Metric. We employ dodging success rate (DSR) as the primary evaluation metric for AFR. DSR is computed using a thresholding strategy for face verification and a closed-set strategy for face identification. For verification, specifically, DSR is defined as:
$\mathrm{DSR}=\frac{1}{N}\sum_{x}\mathbb{I}\big{(}\cos(\mathrm{FR}(x_{\mathrm{p}}),\mathrm{FR}(x_{\mathrm{g}}))>\tau\big{)}\times 100\%,$
where $\mathbb{I}$ is the indicator function, $N$ is the number of face images and $\mathrm{FR}$ represents the target face recognition model. $x_{\mathrm{p}}$ denotes the protected image, while $x_{\mathrm{g}}$ stands for the gallery image of the corresponding identity. The similarity threshold $\tau$ is set at a $0.01$ False Acceptance Rate (FAR) for each victim model. For face identification, we use the Rank-1 Accuracy, which measures whether the top-1 candidate list excludes the original identity of $x$.
To assess the visual quality, we use LPIPS, PSNR, and SSIM [33]. A lower LPIPS score reflects higher perceptual similarity, while higher PSNR and SSIM values indicate better pixel-wise and structural fidelity to the original image.
Implementation Detail. MASQUE builds on the pre-trained Stable Diffusion v1.4 model, using DDIM denoising over $T=50$ steps with a fixed guidance scale of $7.5$. During the backward diffusion process, CA injection occurs in $[T,T-\tau_{\mathrm{attn}}]$ ($\tau_{\mathrm{attn}}=40$), localize optimization in $[T,T-\tau_{\mathrm{edit}}]$ ($\tau_{\mathrm{edit}}=5$), and adversarial guidance in $[T-\tau_{\mathrm{adv}},0]$ ($\tau_{\mathrm{adv}}=45$). We set $\lambda_{\mathrm{CoSi}}=0.1$, $\lambda_{\mathrm{LPIPS}}=1$, and cap optimization to $\mathrm{max}\_\mathrm{adv}=15$ iterations.
5.2 Main Comparison Results
Against Black-Box FR Models. Table 2 presents the DSR for both face verification and Rank-1 identification in black-box scenarios, evaluated using four widely used pre-trained FR feature extractors [6, 11, 25, 2]. For each target model, the remaining three serve as surrogate models to simulate the black-box scenario, with results averaged across three makeup styles. Our method significantly outperforms all baselines, achieving an average DSR of $85.50\%$ for identification and $84.92\%$ for verification. In addition, while MASQUE without a guide image achieves a reasonable DSR, it still falls short compared to its guided counterpart. Furthermore, image quality degrades significantly without a guide image (see supporting evidence in Appendix B.2.).
(a) Face++
(b) Luxand
Figure 4: Performance comparisons against commercial FR APIs. Confidence scores (lower is better) from Face++ and Luxand show that MASQUE provides stronger privacy protection than state-of-the-art noise-based and makeup-based AFR approaches.
Real-world Performance. We also test our method against the commercial APIs, Face++ and Luxand, in verification mode, which assigns similarity scores from 00 to $100$. Figure 4 shows the results. As a proprietary model with unknown training data and parameters, it serves as a realistic testbed for evaluating the effectiveness of AFR methods. Our approach achieves the lowest similarity score across both APIs, demonstrating effectiveness in both open-source and closed-source settings, reinforcing its real-world applicability.
TIP-IM AMT-GAN C2P(I) C2P(D) DiffAM MASQUE
LPIPS (↓) $0.311$ $0.342$ $0.460$ $0.473$ $0.399$ $0.294$
PSNR (↑) $32.158$ $19.512$ $18.922$ $17.990$ $18.314$ $25.815$
SSIM (↑) $0.928$ $0.613$ $0.583$ $0.563$ $0.769$ $0.856$
Table 3: Quantitative comparisons of the visual quality across different AFR methods. Lower values are better for LPIPS, while higher values are better for PSNR and SSIM.
Visual Quality. Our method achieves superior image quality across multiple evaluation metrics, as summarized in Table 3 and illustrated in Figure 1. While TIP-IM attains the highest PSNR and SSIM due to its small perturbation budget, these pixel-level metrics often fail to reflect perceptual quality. In contrast, our approach prioritizes perceptual consistency, balancing content fidelity and visual realism, as demonstrated by its strong LPIPS performance. We also confirm the advantages of MASQUE over existing AFR methods in achieving precise, localized makeup transfer (see additional experiments in Appendix B.1).
Prompt Adherence. We assess prompt adherence in terms of spatial precision and color accuracy. For spatial adherence, we use Concentration, which quantifies the proportion of modifications within a binary mask of the target area. Higher scores indicate better localization, while lower scores suggest spillover. For color precision, we evaluate Hue Accuracy by extracting modified pixels within the mask, converting them to HSV space, and measuring how many fall within the expected hue range. Figure 3 shows that MASQUE consistently achieves the highest concentration scores and competitive hue accuracy across prompts, ensuring precise localization with accurate color application. In contrast, DiffAM struggles with blue eyebrows, likely due to its framework restricting makeup transfer to specific facial regions, limiting its flexibility. These results confirm that MASQUE achieves strong prompt adherence by maintaining spatial precision while accurately reflecting the intended color attributes.
(a) $\tau_{\mathrm{adv}}$
(b) $\mathrm{max}\_\mathrm{adv}$
Identification DSRVerification DSRLPIPS
Figure 5: Ablation studies on hyperparameters $\tau_{\mathrm{adv}}$ and $\mathrm{max}\_\mathrm{adv}$, highlighting the trade-off between DSR and image quality.
5.3 Ablation Studies
Adversarial Strength vs. Image Quality. Increasing $\tau_{adv}$ enhances the adversarial effect, improving DSR but degrading image quality, as indicated by higher LPIPS values in Figure 5. Similarly, increasing $\mathrm{max}\_\mathrm{adv}$ amplifies protection but further compromises image quality.
Customized Makeup. Figure 6 demonstrates the controllability of MASQUE across the following three distinct aspects. Additional visualizations are available in the Appendix C.
Sequential Editing. MASQUE enables progressive makeup application, allowing users to apply makeup sequentially to multiple facial areas. Each additional layer of makeup enhances the protection effect, as reflected in the decreasing similarity scores shown in Figure 6.
Intensity Control. The intensity of the makeup can be adjusted via text prompts. A simple modifier like “vivid” alters the cross-attention, resulting in color changes. Intensity variation has minimal effect on adversarial effectiveness.
Style Control. MASQUE supports full-face transformations. Prompts like “tanned skin” can blend naturally, while dramatic styles like “clown makeup” are also well-adapted. Greater makeup coverage further strengthens protection.
Figure 6: Examples showing the controllability of MASQUE through text prompts. The number at the bottom right indicates the similarity score between the image and the corresponding gallery image.
6 Further Discussions
Guide Images. While MASQUE with guide images enhance DSR and image quality, they may not always be available. As an alternative, we test self-augmented inputs as guide images via flipping and translation. However, preliminary results in Appendix B show limited improvements, indicating the need for more effective strategies to substitute the guide image.
Computational Efficiency. MASQUE enables adversarial makeup applications without pretraining. However, due to the iterative nature of diffusion models, inference speed remains a computational bottleneck. Future work can explore acceleration schemes to improve the efficiency of MASQUE.
Adaptive Evaluations. Real-world AFR threats evolve, with attackers refining strategies over time [7]. Considering adaptive scenarios, such as post-processing transformations (e.g., compression, smoothing, and makeup removal) and adversarial fine-tuning, could help improve robustness and provide a more comprehensive evaluation of AFR methods.
@@ 7 Conclusion
We introduced MASQUE, a diffusion-based AFR framework that applies adversarial makeup via text prompts, enabling localized modifications with automatically generated masks. Unlike prior AFR methods, MASQUE achieves high DSR while preserving visual quality without relying on external IDs, mitigating the potential risks of targeted misuse. Our experiments demonstrated the strong performance of MASQUE, where we highlight the important role of pairwise adversarial guidance. Future work may focus on real-time optimization and robustness against evolving FR models to further strengthen facial privacy protection.
@@
The code implementation of MASQUE and our experiments are open-sourced and available at https://github.com/TrustMLRG/MASQUE.
Appendix A Related Work
A.1 Anti-Facial Recognition
Earlier works on AFR proposed to use obfuscation techniques to obscure the facial identity features [20] or craft $\ell_{p}$-norm bounded perturbations to fool FR models [37]. While effective, these methods often compromise image quality, limiting their practicality in real-world scenarios. Poisoning-based methods [4, 27] introduced a new approach by injecting subtle adversarial noise into images to degrade the effectiveness of recognition models. These techniques excel at disrupting model training or inference without visibly altering the image, but their reliance on model-specific perturbations limits their generalizability across diverse architectures. Unrestricted adversarial examples, which aim to create natural-looking modifications, marked a significant advancement by integrating adversarial signals into realistic alterations without relying on specific constraints. These strategies leverage the flexibility of generative models to balance visual coherence and adversarial effectiveness.
Adversarial Makeup. Adversarial makeup [38, 12, 26, 30] has emerged as a practical solution to realize the goal of facial privacy protection. These methods leverage the strong generative capability of generative models to embed adversarial perturbations into natural makeup-based facial modifications, deceiving attackers’ facial recognition models while largely preserving the aesthetic appeal. For example, [12] proposed AMT-GAN, which introduces a regularization module and a joint training pipeline for adversarial makeup transfer within the generative adversarial network (GAN) framework [8]. Recent advancements in generative models have enhanced the performance of adversarial makeup transfer, such as Clip2Protect [26] and DiffAM [30], which adopt text-guided StyleGAN model and diffusion-based framework, respectively, to enable seamless adversarial face modifications with much improved visual quality. In this work, we build upon these developments by leveraging diffusion models to generate visually consistent, localized adversarial makeup for privacy protection in dodging scenarios.
A.2 Diffusion-Based Image Editing
Localized Image Editing. Localized editing methods extend diffusion models for targeted modifications. For example, mask-based approaches [1, 5] constrain edits to specific regions using spatial masks, preserving unedited areas but often struggling with structural consistency. Attention-based approaches [9, 32] guide edits via attention injection, achieving better global structure preservation but suffering from unintended changes (editing leakage). Our proposed method, MASQUE, utilizes both mask-based and attention-based guidance without any fine-tuning, combining mask precision with attention flexibility, ensuring semantically consistent and region-specific modifications while addressing the limitations of prior approaches.
Text-Guided Diffusion Models. Text-guided diffusion models are a major advancement in generative AI, capable of synthesizing high-quality images from natural language descriptions. Built on denoising diffusion probabilistic models (DDPMs) [10], they iteratively refine random noise into coherent images guided by text prompts. Recent advancements in diffusion models have expanded their capabilities to tasks such as localized editing and controllable generation. Mask-based approaches [1, avrahami2022blended, 5] achieve local text-guided modifications by incorporating user-defined constraints, like masks, to confine edits to specific regions. While effective in preserving unedited areas, these methods often struggle to maintain structural consistency within the edited regions, especially in complex scenarios. On the other hand, mask-free attention-based methods [9, cao2023masactrl, 32] use attention injection mechanisms to guide edits without requiring explicit masks. These methods excel at preserving the global structure of the image but are prone to editing leakage, where changes unintentionally affect areas beyond the intended region. In this work, we leverage mask-based cross-attention guidance [18] to achieve localized adversarial perturbations in the form of makeup. This approach ensures the precise application of modifications to desired regions, guided by user-specified prompts while addressing the limitations of structural inconsistencies seen in previous methods.
A.3 Makeup Transfer
Advancements in makeup transfer have moved from GAN-based methods to more advanced diffusion models. Traditional approaches like BeautyGAN [16] and PSGAN [17] use histogram matching, attention mechanisms, and spatial encodings to transfer makeup while preserving structure. CPM [21] expands beyond color transfer using UV mapping, while RamGAN [35] ensure component consistency. However, GAN-based models struggle with extreme styles and rely on imprecise pseudo-paired data, limiting fidelity.
To overcome these challenges, diffusion-based methods like SHMT [31] and Stable-Makeup [40] offer improved realism and robustness. Specifically, SHMT eliminates reliance on pseudo-paired data through a self-supervised “decoupling-and-reconstruction” framework, leveraging a Laplacian pyramid for hierarchical texture transfer. Stable-Makeup employs a diffusion model with a Detail-Preserving encoder and cross-attention layers to ensure precise, structure-preserving makeup application. These approaches set a new standard for high-fidelity transfer. However, our study takes a different approach—rather than transferring makeup from a reference image, we focus on text-guided makeup generation using diffusion models, enabling intuitive and flexible image editing to create customizable makeup based solely on textual descriptions. This also eliminates the need for model fine-tuning on specific makeup images, allowing us to generate the desired makeup directly during inference, resulting in a more efficient and adaptable approach.
Appendix B Additional Experiments
B.1 Localized Editing
To evaluate whether edits are confined to the desired region, images are divided into in-mask and out-mask regions using a binary mask, isolating the corresponding areas in both images via element-wise multiplication. The areas are then used to normalize similarity metrics proportionally, ensuring fair comparisons. Specifically, we employ DISTS, LPIPS, and PieAPP as evaluation metrics in this experiment: DISTS measures perceptual dissimilarity based on structure and texture, LPIPS uses deep features, and PieAPP reflects human perceptual preferences.
Metric Method In-Mask Out-Mask $\textbf{Diff}\ \bm{\Delta}$
DISTS ()($\downarrow$)( ↓ ) TIP-IM $2.013$ $0.094$ $1.919$
AMT-GAN $7.993$ $0.149$ $7.844$
C2P (I) $8.945$ $0.153$ $8.792$
C2P (D) $10.011$ $0.169$ $9.842$
DiffAM $9.645$ $0.141$ $9.504$
MASQUE $12.306$ $0.104$ $12.202$
LPIPS ()($\downarrow$)( ↓ ) TIP-IM $0.322$ $0.305$ $0.017$
AMT-GAN $0.360$ $0.316$ $0.044$
C2P (I) $0.400$ $0.447$ $-0.047$
C2P (D) $0.436$ $0.460$ $-0.025$
DiffAM $0.509$ $0.410$ $0.099$
MASQUE $0.434$ $0.290$ $0.144$
PieAPP ()($\downarrow$)( ↓ ) TIP-IM $3.621$ $0.462$ $3.1591$
AMT-GAN $21.325$ $0.814$ $20.511$
C2P (I) $17.186$ $1.138$ $16.049$
C2P (D) $20.850$ $1.314$ $19.536$
DiffAM $17.523$ $1.312$ $16.211$
MASQUE $21.433$ $0.645$ $20.789$
Table 4: Comparison of perceptual similarity metrics across in-mask and out-mask regions and different AFR methods.
Table 4 compares the effectiveness of localized edits, where higher $\Delta$ values indicate stronger localization of perturbations. TIP-IM [37], as a noise-based method, applies pixel-wise adversarial perturbations uniformly, resulting in the smallest difference for all metrics due to minimal distinction between in-mask and out-mask regions. In contrast, our method introduces significant perturbations in the in-mask region, leading to poorer metrics there, but achieves the best or second-best results in the out-mask region, indicating minimal disruption to untouched areas.
B.2 Pairwise Adversarial Guidance
Leveraging multiple guide images further stabilizes adversarial training, reducing bias toward any single representation and improving robustness. To evaluate its effectiveness, we use FaceNet as the target model and select $100$ identities from CelebA-HQ, varying the number of guide images per identity. These evaluation identities are distinct from the main experiments, ensuring at least $12$ images per identity to accommodate setups with up to $10$ guide images.
Effectiveness is measured using the confidence score, the gap between the similarity score of the original identity’s gallery image and that of the misidentified identity. Table 5 shows that increasing guide images enhances identity obfuscation, leading to higher confidence scores while reducing similarity with the original identity. Moreover, images generated with guide images in the pairwise adversarial guidance step exhibit higher visual quality than those relying solely on adversarial loss on the original image. This confirms that simply maximizing distance from the original image is ineffective, as it disrupts the diffusion process’s goal of preserving natural image characteristics, leading to instability in both image quality and adversarial performance. These results demonstrate that MASQUE’s pairwise adversarial guidance provides a robust, privacy-centric solution, balancing strong identity obfuscation with high visual fidelity.
#Guide Iden. DSR()($\uparrow$)( ↑ ) Veri. DSR()($\uparrow$)( ↑ ) Confidence()($\uparrow$)( ↑ ) Similarity()($\downarrow$)( ↓ ) LPIPS()($\downarrow$)( ↓ )
00 $0.66$ $0.60$ $0.114$ $0.355$ $0.303$
$1$ $0.67$ $0.63$ $0.135$ $0.349$ $0.292$
$2$ $0.72$ $0.71$ $0.165$ $0.297$ $0.291$
$5$ $0.74$ $0.74$ $0.177$ $0.283$ $0.291$
$10$ $0.76$ $0.77$ $0.197$ $0.266$ $0.291$
Guide Method Iden. DSR()($\uparrow$)( ↑ ) Veri. DSR()($\uparrow$)( ↑ ) PSNR()($\uparrow$)( ↑ )
Pairwise ($G=1$) $0.72$ $0.66$ $25.691$
Self-Guidance ($G=0$) $0.56$ $0.40$ $25.812$
    + Horizontal Flip $0.57$ $0.50$ $25.834$
    + Vertical Flip $0.12$ $0.06$ $26.566$
    + Translation $0.55$ $0.38$ $25.807$
Table 6: Impact of adversarial image guidance in DSR and PSNR. We test pairwise- and self-guidance with different augmentations.
B.3 Self-Augmented Guidance
To address cases where users may lack suitable guide images, we tested using self-augmented input images as guide images. This experiment was conducted on $100$ identities from CelebA-HQ with the prompt “red lipstick”, targeting the FaceNet model. For augmentation, we applied horizontal flipping, vertical flipping, and random translation with a maximum shift of $20$ pixels. Results, as shown in Table 6, compare these approaches against the input image itself as the guide image. Horizontal flipping and random translation perform similarly to using the input image as the guide, whereas vertical flipping shows minimal adversarial effect. This is likely because vertically flipped faces are hard to align with facial images by the recognition model.
B.4 DSR under Rank-5 Accuracy
AMT-GAN DiffAM C2P (I) C2P (D) MASQUE
Rank-1 $54.39$ $61.25$ $36.92$ $71.17$ $92.08$
Rank-5 $28.58$ $45.58$ $18.25$ $48.67$ $81.42$
Table 7: DSR with Rank-1 and Rank-5 metrics across AFR methods
B.5 Mask and Perturbation Localization
Figure 7: Visualizations of MASQUE with and without masking. The first row uses prompt “shimmery gold eyeshadow”, while the second row uses “blue eyeshadow”.
To understand the impact of masking in MASQUE, we conduct an ablation study with and without applying a mask. Figure 7 shows that the absence of a mask causes the model to struggle with spatial control, leading to unconstrained perturbations across the entire image rather than confining makeup to the intended area.
Appendix C Additional Visualizations
Figure 8: Visualizations of MASQUE-generated images with or without using the edit loss.
Additional visual results of image-guided makeup transfer are provided in Figure 8 to accommodate space limitations in the main text. MASQUE demonstrates precise makeup application for each prompt, regardless of the targeted region or the subject’s demographics.
Appendix D Algorithm Pseudocode
Input: $\bm{z}_{\mathrm{edit}},\bm{z}_{\mathrm{rec}}$: Edited and original latent, $\bm{e}_{\mathrm{edit}},\bm{e}_{\mathrm{rec}}$: Edited and original text embeddings, $\mathcal{D}$: Stable Diffusion model, $T$: Number of diffusion steps, $\tau_{\mathrm{attn}},\tau_{\mathrm{edit}},\tau_{\mathrm{adv}}$: Step thresholds, $\mathcal{M}$: Binary region mask, $\mathcal{I}_{\mathrm{target}}$: Target token indices, $d_{k}$: Key/Query dimensionality, $\lambda_{\mathrm{edit}},\lambda_{\mathrm{CoSi}},\lambda_{\mathrm{LPIPS}}$: Loss weights, $\eta$: Learning rate, $\sigma(t)$: Noise scale function, $\bm{x}$: Input image, $\tilde{\bm{x}}$: Guide image, $max\_adv$: Maximum adversarial iterations

### v2

@@ 5.1 Experimental Setup
Dataset. To reflect the high-quality nature of online face images, we use 300300300 high-resolution images ($\times 1024$ 1024) from each of the CelebA-HQ (Karras et al., 2017) and VGG-Face2-HQ (Chen et al., 2024) datasets. For each dataset, we randomly sample 100 identities, each with three images: a probe $\bm{x}$ (to be protected), a reference $\tilde{\bm{x}}$ (as the guide image), and a gallery image $\bm{x}_{\mathrm{g}}$ (stored in the attacker’s facial recognition model). Note that the VGG-Face2-HQ dataset contains face images under challenging conditions, such as pose variation and natural occlusions like sunglasses and hair.
@@
Configuration. We compare MASQUE against the following baselines: TIP-IM (Yang et al., 2021), AMT-GAN (Hu et al., 2022), C2P (Shamshad et al., 2023), and DiffAM (Sun et al., 2024). For C2P, both its impersonating version, C2P (I), and its dodging version, C2P (D), are assessed. We compare the performance of MASQUE with existing AFR techniques on four public FR models: IR152 (Deng et al., 2019), IRSE50 (Hu et al., 2018), FaceNet (Schroff et al., 2015), and MobileFace (Chen et al., 2018). Without explicitly mentioning, we set $G=1$ in MASQUE (using a single separate guide image of the same individual). For reference-based methods, we selected images from the makeup dataset (Li et al., 2018) used during their pre-training that best matched the given prompt. In addition to public FR models, we evaluate AFR methods against two commercial APIs: Face++ [^1] https://www.faceplusplus.com/face-comparing/ and Luxand [^2] https://luxand.cloud/face-api.
Evaluation Metric. We employ DSR to evaluate the protection effectiveness of AFR methods. We measure visual quality using LPIPS, PSNR, and SSIM (Wang et al., 2004). Lower LPIPS and higher PSNR/SSIM indicate better preservation in terms of perceptual and structural similarity. Besides, we assess prompt following ability using metrics like concentration and hue accuracy. In Appendix B, we provide detailed descriptions of the hyperparameters used in MASQUE, implementation details for external image references, and formal definitions of the evaluation metrics.
5.2 Main Results
(a) Face++, CelebA
(b) Luxand, CelebA
(c) Face++, VGG-Face2
(d) Luxand, VGG-Face2
Figure 3: Performance comparisons between different AFR methods against two commercial APIs.
[…]
Table 3: Visual quality of various AFR methods on CelebA-HQ averaged across three makeup styles.
TIP-IM AMT-GAN C2P(I) C2P(D) DiffAM MASQUE ($G=0$) MASQUE ($G=1$)
LPIPS($\downarrow$) 0.3110.3110.311 0.3420.3420.342 0.4600.4600.460 0.4730.4730.473 0.3990.3990.399 0.2980.2980.298 0.2940.2940.294
PSNR($\uparrow$) 32.1632.1632.16 19.5119.5119.51 18.9218.9218.92 17.9917.9917.99 18.3118.3118.31 25.3325.3325.33 25.8225.8225.82
SSIM($\uparrow$) 0.9280.9280.928 0.6130.6130.613 0.5830.5830.583 0.5630.5630.563 0.7690.7690.769 0.8390.8390.839 0.8560.8560.856
Visual Quality. MASQUE achieves superior image quality across multiple evaluation metrics, as summarized in Table 3. While TIP-IM attains the highest PSNR and SSIM due to its small perturbation constraint, these pixel-level metrics often fail to reflect perceptual quality (see Figure 1 for qualitative comparison results). In contrast, our approach prioritizes perceptual consistency, balancing content fidelity and visual realism, as demonstrated by its strong LPIPS performance.
Figure 4: Prompt adherence of MASQUE. Similarity score is shown at the bottom right.
Moreover, we examine the localization capability and prompt adherence of MASQUE based on spatial precision and color accuracy. For spatial adherence, we use concentration, which quantifies the proportion of modifications within a binary mask of the target area. Higher scores indicate better localization, while lower scores suggest spillover. For color precision, we evaluate hue accuracy by extracting modified pixels within the mask, converting them to HSV space, and measuring how many fall within the expected hue range. Figure 5 shows that MASQUE consistently achieves the highest concentration scores and competitive hue accuracy across prompts, ensuring precise localization with accurate color application. In contrast, DiffAM struggles with blue eyebrows, likely due to its framework restricting makeup transfer to specific facial regions, limiting its flexibility. In Appendix D.1, we conduct additional experiments using other distance metrics with respect to both in- and out-mask regions, illustrating the strong localization capability of MASQUE. These results confirm that MASQUE achieves strong prompt adherence by maintaining spatial precision while accurately reflecting the intended color attributes.
@@ 5.2 Main Results
Figure 5: Comparison of concentration and hue accuracy on CelebA-HQ across different AFR methods for three makeup styles. Concentration values are normalized relative to MASQUE.
6 Further Analysis
Controllable Makeup Generation. Figure 4 illustrates the strong controllability of MASQUE over text makeup prompts across three dimensions: sequential editing, intensity control, and style control. Makeup can be applied progressively to multiple facial regions, with each layer enhancing protection, as shown by decreasing similarity scores. Intensity is adjustable via text prompts (e.g., “vivid”), affecting appearance with minimal impact on effectiveness. Style prompts range from natural (e.g., “tanned skin”) to extreme (e.g., “clown makeup”), with broader coverage offering stronger protection (see Appendix E for additional visualizations of MASQUE-generated images).
Table 4: DSR (%) of MASQUE on CelebA-HQ and “red lipstick” against various image transformations.
Method Identification Verification
IR152 IRSE50 MobileFace FaceNet IR152 IRSE50 MobileFace FaceNet
MASQUE 1.001.001.00 0.970.970.97 0.940.940.94 0.710.710.71 0.970.970.97 0.950.950.95 0.690.690.69 0.660.660.66
  + Resize 0.970.970.97 0.950.950.95 0.910.910.91 0.710.710.71 0.920.920.92 0.920.920.92 0.670.670.67 0.650.650.65
  + Compression 1.001.001.00 0.950.950.95 0.920.920.92 0.720.720.72 0.960.960.96 0.930.930.93 0.650.650.65 0.630.630.63
  + Gaussian Noise 0.980.980.98 0.950.950.95 0.950.950.95 0.820.820.82 0.960.960.96 0.920.920.92 0.900.900.90 0.810.810.81
  + Blur 1.001.001.00 0.950.950.95 0.910.910.91 0.730.730.73 0.970.970.97 0.930.930.93 0.700.700.70 0.650.650.65
  + Fog 0.950.950.95 0.970.970.97 0.930.930.93 0.880.880.88 0.920.920.92 0.970.970.97 0.840.840.84 0.810.810.81
  + Makeup Removal 0.950.950.95 0.940.940.94 0.910.910.91 0.850.850.85 0.930.930.93 0.920.920.92 0.690.690.69 0.500.500.50
  + DDPM Denoising 0.880.880.88 0.880.880.88 0.770.770.77 0.670.670.67 0.750.750.75 0.840.840.84 0.850.850.85 0.670.670.67
Robustness to Image Transformation. We evaluate the robustness of MASQUE using the same 100100100 CelebA-HQ images and four FR models used in the main experiments, with the prompt “red lipstick”. We then subject the edited images to a variety of transformations, standard post-processing such as JPEG compression, blurring, natural corruptions like ImageNet-C, and adaptive attacks such as DiffAM’s makeup removal and DDPM-based denoising. As shown in Table 4, MASQUE maintains high dodging success rates across all cases, indicating strong resilience to real-world degradations and adaptive threats. This robustness may arise from the invariance of FR models, which are trained to withstand such transformations, allowing MASQUE’s protection to persist. Also, we test the performance of MASQUE against adversarial purification (Nie et al., 2022), where the results are presented and discussed in Appendix D.4.
Computational Analysis. MASQUE does not require fine-tuning, operating directly with a pretrained Stable Diffusion model. In contrast, Clip2Protect fine-tunes per image (~252525s), and DiffAM performs per-dataset fine-tuning (~373737min) for each style or identity. At inference, MASQUE runs a one-time iterative optimization (~909090s), slightly slower than Clip2Protect (~181818s) and DiffAM (~303030s). Despite this, MASQUE offers key advantages: it supports arbitrary prompts and identities without retraining, making it more flexible and general. Under typical AFR scenarios, where protection is applied once before sharing, this adaptability outweighs the modest runtime difference.
@@ 7 Conclusion
We introduced MASQUE, a diffusion-based anti-facial recognition framework that applies adversarial makeup via text prompts, enabling localized modifications with automatically generated masks. Unlike prior AFR methods, MASQUE achieves high dodging success rates while preserving visual quality without relying on external IDs, mitigating the potential risks of targeted misuse. Our experiments demonstrated the superiority of MASQUE over existing baselines, where we highlight the important role of pairwise adversarial guidance. Future work may focus on optimizing the efficiency for real-time applications and studying the robustness against adaptively evolving FR models (Fan et al., 2025) to further strengthen MASQUE’s efficacy for facial privacy protection.
@@
The code implementations of our method and for reproducing our experimental results are open-sourced and available at: https://github.com/TrustMLRG/MASQUE.
Appendix A Algorithm Pseudocode
Algorithm 1 presents the pseudocode of MASQUE, which consists of modules such as precise null-text inversion, cross-attention fusion with masking, and pairwise adversarial guidance.
Input: $\bm{z}_{\mathrm{edit}},\bm{z}_{\mathrm{rec}}$: edited and original latent, $\bm{e}_{\mathrm{edit}},\bm{e}_{\mathrm{rec}}$: edited and original text embeddings, $\mathcal{D}$: Stable Diffusion model, $T$: number of diffusion steps, $\tau_{\mathrm{attn}},\tau_{\mathrm{edit}},\tau_{\mathrm{adv}}$: step thresholds, $\mathcal{M}$: binary region mask, $\mathcal{I}_{\mathrm{target}}$: target token indices, $d_{k}$: key/query dimensionality, $\lambda_{\mathrm{edit}},\lambda_{\mathrm{CoSi}},\lambda_{\mathrm{LPIPS}}$: loss weights, $\eta$: learning rate, $\sigma(t)$: noise scale function, $\bm{x}$: input image, $\tilde{\bm{x}}$: guide image, $m_{\mathrm{adv}}$: maximum adversarial iterations